You mix 258 ml of 0.613 M CsoH with 510.0 mL of 0.3101 M HF in a coffee cup calorimeter, and the temperature of both solutions rises fron 20.50°C before mixing, to 23.50°C after the reaction. CsoH (aq) HF (aq) CsF (aq) H20 (L) What is the heat [q] for this reaction per mole of CsoH? Assume the densities of the solutions are all 1.00 g/mL and the specific heats of the solutions are 4.18 J/(g. K) 93.5 kJ/mol 60.9 kJ/mol 71.3 kJ/mol 55.9 kJ/mol 74.4 kJ/mol
Answer #1

mol of CsOH reacting = M*V

= 0.613 M * 0.258 L

= 0.158 mol

mol of HF reacting = M*V

= 0.3101 M * 0.510 L

= 0.158 mol

total volume of solution = 258 mL + 510 mL

= 768 mL

since density of solution is 1 g/mL,

mass of solution = 768 g

Q solution = m*C*delta T

= 768 * 4.18 * (23.50 - 20.50)

= 9630.7 J

= 9.631 KJ

q per mol = 9.631 KJ/0.158 mol

= 60.9 KJ/mol

Answer: 60.9 KJ/mol